16x2+4y2+32x-16y-32=0

center,foci, ends of major and minor axis, ends of latus rectum


1
Expert's answer
2021-08-02T16:46:26-0400

Consider the equation "16x^2+4y^2+32x-16y-32=0"

"16(x^2+2x)+4(y^2-4y)=32"

"16(x^2+2x+1-1)+4(y^2-4y+4-4)=32"

"16(x^2+2x+1)-16+4(y^2-4y+4)-16=32"

"16(x+1)^2+4(y-2)^2=32+16+16"

"16(x+1)^2+4(y-2)^2=64"

Dividing by 64, we get

"\frac{(x+1)^2}{4}+\frac{(y-2)^2}{16}=1"

Observe that the above equation is of the form

"\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1"

So, the given equation is ellipse

Center of the ellipse "=(h,k)=(-1,2)"

And "a^2=4,b^2=16"

Observe that "a^2<b^2"

To find the foci of the ellipse, find "c" using the relation "c^2=b^2-a^2"

"c^2=16-4=12"

"\Rightarrow c=\pm 2\sqrt{3}"

Foci of the ellipse are: "(h,k-\sqrt{c}),(h,k+\sqrt{c})=(-1,2-2\sqrt{3}),(-1,2+2\sqrt{3})"

Ends of the major axis are: "(h,k-b),(h,k+b)=(-1,2-4),(-1,2+4)=(-1,-2),(-1,6)"

Ends of the minor axis are:

"(h-a,k),(h+a,k)=(-1-2,2),(-1+2,2)=(-3,2),(1,2)"

Ends of the latus rectum are:

"(\frac{-a^2}{b},c),(\frac{a^2}{b},c)=(-1,2\sqrt{3}),(1,2\sqrt{3})"